# What no one has seen before: gravitational waveforms from warp drive collapse

### v1

@@ 3.2 Matter fluxes
A visualisation of the evolution of the matter-energy content is shown in Fig. 3. In Fig. 5, we show a comparison of the matter and gravitational-wave fluxes out of the sphere at a coordinate radius of $r=14[R]$ around the warp bubble versus time. We see that whilst the gravitational-wave flux is strictly positive as expected, the matter flux alternates in sign, with waves of positive and negative energy leaving the volume. It appears from the evolution that the ejection of matter is driven by a shell of positive energy forming within the negative energy density shell of the initial bubble and driving it outwards. This is reminiscent of the early works of Bondi [42], showing the effect of repulsion and attraction between positive and negative energy matter.
In Fig. 6, we show a zoom-in of the fluxes and integrated energy loss, which shows that the matter flux results in an overall negative flux out of the spacetime, with the gravitational waves giving a much smaller positive flux. The result is that the overall ADM mass of the spacetime will become more positive (having started at zero). It would be interesting to study whether this is a generic effect, such that the ADM energy for spacetimes with NEC-violating matter does not dynamically decrease below zero. However, we can see from Fig. 6 that nothing appears to impede the gravitational-wave flux when the net matter flux is instantaneously positive overall, as between times $t-r_{ext}=-1.0$ and 0.5-0.5 on the left panel of the plot.
@@ 4 Discussion
We also studied the evolution of the fluxes of matter and gravitational-wave energy from the spacetime, finding that an initial wave of negative energy matter is emitted, followed by alternating positive and negative energy waves. The gravitational-wave flux starts shortly after and is positive throughout, as expected. The end result is a net negative flux of energy, resulting in an increased ADM mass of the spacetime. It would be interesting to explore to what extent this behaviour is generic or depends on the chosen equation of state supporting the warp bubble. To do this, one would want to consistently implement a fluid description with a general rest frame velocity and equation of state, unlike the somewhat ad hoc setup employed here.
This work could be extended in several other directions. Firstly, using a code better adapted to the cylindrical symmetry, one could extend to higher speeds, including those that exceed $v=1.0$, to see if there is a change in behaviour as light speed is approached. One might expect to see some kind of pathology develop, e.g., the formation of elliptic regions that cause the dynamical evolution to fail. Better resolution and longer evolutions would also allow better characterization of the strain signal, particularly the role of the gravitational-wave memory content (see [65]).
As discussed above, for a 1km-sized ship, the frequency of the signal is much higher than the range probed by existing detectors, and so current observations cannot constrain the occurrence of such events. However, the amplitude of the strain signal would be significant for any such event within our galaxy and even beyond, and so within the reach of future detectors targeting higher frequencies [60]. We caution that the waveforms obtained are likely to be highly specific to the model employed, which has several known theoretical problems, as discussed in the Introduction. Further work would be required to understand how generic the signatures are, and properly characterise their detectability.
@@ 5 Acknowledgements
The crew thanks the creators and actors of the Star Trek universe for their inspiration for this work. We also thank the many people who discussed the project and gave useful feedback and encouragement, especially the 2023 Peyresque workshop attendees and our NR friends at the 2023 ETK meeting in Aveiro. Particular conceptual help was given by Jean Alexandre, Josu Aurrekoetxea, Juan Valiente Kroon, Gautam Satischandran, and Miren Radia and Josu Aurrekoetxea provided the gravitational-wave analysis scripts used. The GRTL collaboration is warmly acknowledged for the use of their GRChombo code (www.grchombo.org) and the development work that supports it. Katy “I cannae change the laws of physics Captain” Clough thanks the University of Potsdam and the Albert Einstein Institute in Potsdam for hosting her for the final stage of the project and acknowledges current support from an STFC Ernest Rutherford fellowship, project reference ST/V003240/1, and previous support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 693024). Tim “Captain Jim T.” Dietrich thanks Pedro Ferreira for hosting him at the University of Oxford, which allowed this project to make significant progress. TD also acknowledges support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Sebastian “Wrath of” Khan thanks Jens Niemeyer for hosting him in Georg August University of Goettingen, where the voyage towards this article first began. The simulations presented in this paper used the Glamdring cluster in Astrophysics at Oxford University and DiRAC resources under the projects of dp092 (Testing the limits of General Relativity). This work used the DiRAC Memory Intensive service (Cosma8 and Cosma7) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. This work also used the DiRAC Data Intensive service (DIaL3) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. This work also used the DiRAC Data Intensive service (CSD3) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.
@@ Appendix A Numerical details, validation and convergence
$\mathcal{M}_{i}$ $:=\tilde{\gamma}^{kl}\left(\partial_{k}\tilde{A}_{li}-2\tilde{\Gamma}^{m}_{l(i}\tilde{A}_{k)m}-3\tilde{A}_{ik}\frac{\partial_{l}\chi}{2\chi}\right)-\frac{2}{3}\partial_{i}K-8\pi S_{i}.$ (25)
@@
To verify that the matter is sufficiently resolved within the volume of interest, we also reconcile the matter quantities over time, checking that the net energy flux out of the surface matches the change in the matter energy and any exchange with the curvature (often referred to as a “source”) over time. For this, we follow the approach of [66]. As shown in the last panel of Fig. 7, the agreement is good throughout the simulation, which gives us confidence in the matter evolution.

### v2

@@ 3.2 Matter fluxes
We measure the matter flux out of a spherical coordinate surface $\partial\Sigma$ with unit coordinate normal vector $s^{i}$ and coordinate area element $dA$ as
$F=\int_{\partial\Sigma}\sqrt{\gamma}~T^{i}_{0}s_{i}dA$ (24)
where $\gamma$ is the determinant of the spatial metric $\gamma_{ij}$ (see Clough (2021) for a derivation and details on practical implementation). We can also integrate this to find the total energy flux that is emitted up to a time $t$ as
$\Delta E=\int_{0}^{t}F(\tilde{t})d\tilde{t}~.$ (25)
A visualisation of the evolution of the matter-energy density $\rho$ is shown in Fig. 3. In Fig. 5, we show a comparison of the integrated matter and gravitational-wave fluxes out of the sphere at a coordinate radius of $r=14[R]$ around the warp bubble versus time. We see that whilst the gravitational-wave flux is strictly positive as expected, the matter flux alternates in sign, with waves of positive and negative energy leaving the volume. It appears from the evolution that the ejection of matter is driven by a shell of positive energy forming within the negative energy density shell of the initial bubble and driving it outwards. This is reminiscent of the early works of Bondi (1957), showing the effect of repulsion and attraction between positive and negative energy matter.
In Fig. 6, we show a zoom-in of the fluxes and integrated energy loss, which shows that the matter flux results in an overall negative flux out of the spacetime, with the gravitational waves giving a much smaller positive flux. The result is that the final mass of the spacetime volume is more positive (having started at zero). It would be interesting to study whether this is a generic effect, such that spacetimes with NEC-violating matter cannot dynamically settle into a state with a mass below zero. We can see from Fig. 6 that nothing appears to impede the gravitational-wave flux when the net matter flux is instantaneously positive overall, as between times $t-r_{ext}=-1.0$ and 0.5-0.5 on the second panel of the plot so temporary decreases appear to be permitted.
@@
The evolution does appear to settle down into a stable remnant, that could be described as a “star” of the warp fluid - it is a configuration of the fluid that appears to balance the gravitational forces and fluid pressures. We would need to evolve for longer and use a larger grid (to avoid boundary reflection effects) to be sure about its stability, but it appears to have settled to at least an approximately steady state in the period simulated.
Figure 5: A comparison of the integrated matter and gravitational-wave energy fluxes out of sphere at a coordinate radius of $r=14[R]$ around the warp bubble versus time for different $v$. The top panel shows the integrated matter fluxes (i.e., the total energy radiated). The bottom panel shows the integrated gravitational wave fluxes in the 20 mode. The net gravitational-wave flux is always positive, whereas the net matter flux oscillates as waves of positive and negative energy leave the volume. Note that the matter flux is significantly larger in magnitude than that of the gravitational waves.
Figure 6: Zoom in on the start and end periods of the integrated energy fluxes in the matter and gravitational waves as in Fig. 5 for the case $v=0.1$. We see that the flux begins in the matter sector, with alternating waves of positive and negative energy. The gravitational-wave flux starts a little later but continues throughout the periods in which the net matter flux is both positive and negative. The second panel shows that the overall matter energy loss is negative (i.e. the energy of matter in the volume increases), whereas the gravitational-wave flux is overall positive as expected but much smaller. The final mass of the spacetime volume once the spacetime settles into a stationary state
is therefore larger than its initial value (which was zero).
@@ 4 Discussion
We also studied the evolution of the fluxes of matter and gravitational-wave energy from the spacetime, finding that an initial wave of negative energy matter is emitted, followed by alternating positive and negative energy waves. The gravitational-wave flux starts shortly after and is positive throughout, as expected. The end result is a net negative flux of energy, resulting in a higher final mass of the spacetime volume. It would be interesting to explore to what extent this behaviour is generic or depends on the chosen equation of state supporting the warp bubble. To do this, one would want to consistently implement a fluid description with a general rest frame velocity and equation of state, unlike the somewhat ad hoc setup employed here.
This work could be extended in several other directions. Firstly, using a code better adapted to the axisymmetry, one could extend to higher speeds, including those that exceed $v=1.0$, to see if there is a change in behaviour as light speed is approached. One might expect to see some kind of pathology develop, e.g., the formation of elliptic regions that cause the dynamical evolution to fail. Better resolution and longer evolutions would also allow better characterization of the strain signal, particularly the role of the gravitational-wave memory content (Mitman et al., 2024).
As discussed above, for a 1km-sized ship, the frequency of the signal $\sim 300$ kHz is much higher than the range probed by existing detectors, so current observations cannot constrain the occurrence of such events. However, the amplitude of the strain signal would be significant for any such event within our galaxy and even beyond - at a distance of 1 Mpc, slightly further than the Andromeda galaxy, the strain amplitude is similar to LIGO detector’s peak sensitivity. The signal is therefore potentially within the reach of future detectors targeting higher frequencies (Aggarwal et al., 2021). We caution that the waveforms obtained are likely to be highly specific to the model employed, which has several known theoretical problems, as discussed in the Introduction. Further work would be required to understand how generic the signatures are, and properly characterise their detectability.
@@ 5 Acknowledgements
The crew thanks the creators and actors of the Star Trek universe for their inspiration for this work. We also thank the many people who discussed the project and gave useful feedback and encouragement, especially the 2023 Peyresque workshop attendees and our NR friends at the 2023 ETK meeting in Aveiro. Particular conceptual help was given by Jean Alexandre, Josu Aurrekoetxea, Mark Hannam, Juan Valiente Kroon, Gautam Satischandran, and Miren Radia and Josu Aurrekoetxea provided the gravitational-wave analysis scripts used. The GRTL collaboration is warmly acknowledged for the use of their GRChombo code (www.grchombo.org) and the development work that supports it. Katy “I cannae change the laws of physics Captain” Clough thanks the University of Potsdam and the Albert Einstein Institute in Potsdam for hosting her for the final stage of the project and acknowledges current support from an STFC Ernest Rutherford fellowship, project reference ST/V003240/1, and previous support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 693024). Tim “Captain Jim T.” Dietrich thanks Pedro Ferreira for hosting him at the University of Oxford, which allowed this project to make significant progress. TD also acknowledged support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Sebastian “Wrath of” Khan thanks Jens Niemeyer for hosting him in Georg August University of Goettingen, where the voyage towards this article first began.
The simulations presented in this paper used the Glamdring cluster in Astrophysics at Oxford University and DiRAC resources under the projects of dp092 (Testing the limits of General Relativity). This work used the DiRAC Memory Intensive service (Cosma8 and Cosma7) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. This work also used the DiRAC Data Intensive service (DIaL3) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. This work also used the DiRAC Data Intensive service (CSD3) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.
@@ Appendix A Numerical details, validation and convergence
$\mathcal{M}_{i}$ $:=\tilde{\gamma}^{kl}\left(\partial_{k}\tilde{A}_{li}-2\tilde{\Gamma}^{m}_{l(i}\tilde{A}_{k)m}-3\tilde{A}_{ik}\frac{\partial_{l}\chi}{2\chi}\right)$
$\quad-\frac{2}{3}\partial_{i}K-8\pi S_{i}.$ (27)
@@
To verify that the matter is sufficiently resolved within the volume of interest, we also reconcile the matter quantities over time, checking that the net energy flux out of the surface matches the change in the matter energy and any exchange with the curvature (often referred to as a “source”) over time. For this, we follow the approach of Clough (2021). As shown in the last panel of Fig. 7, the agreement is good throughout the simulation, which gives us confidence in the matter evolution.